OpenStudy (anonymous):

Linear Equation in 2 unknown. solve for x+y using 4 methods. A. Graphical Method. B. By Elimination. C. By Substitution D. Determinant.

OpenStudy (anonymous):

Equation: 3x-2y=4 4x-3y=5

OpenStudy (anonymous):

@Broskishelleh here.

OpenStudy (broskishelleh):

I'll try

OpenStudy (broskishelleh):

A.(2,1) Here is the graph

OpenStudy (anonymous):

@Broskishelleh THE 3 METHODS? B. C. D.

OpenStudy (anonymous):

ok, i got the B. now (3x-2y=4) 4 = 12x -8y = 16 (4x-3y=5) 3 = 12x -9y = 17 ---- change the sign. ________________ y = 1 subtitute. 3x-2(1)= 4 3x= 2+4 3x = 6 x= 2

OpenStudy (anonymous):

i need C. and D.

OpenStudy (anonymous):

ok, i got all.